# Spatial Proportional Hazards Model with Differential Regularization

### v1

@@
Spatial Proportional Hazards Model with Differential Regularization
Lorenzo Tedesco
@@ 1 Introduction
Survival analysis has long been an essential tool across various fields, including medicine, engineering, and social sciences, for analyzing time-to-event data. In this paper, we focus on one of the most widely used models within survival analysis: the Proportional Hazards model, first introduced by Cox (1972). This model assesses the effect of covariates on the hazard, or risk, of an event occurring.
@@
In recent years, survival analysis has made significant strides in modeling different forms of dependence in survival data. This paper explores a scenario where the occurrence of an event is associated with its geographic location, introducing a spatial dependence into the data. Specifically, an event’s occurrence may be influenced by spatial factors related to the location where the event happens. We refer to the Proportional Hazards model that incorporates these spatial factors as the Spatial Proportional Hazards model.
Spatial factors can generally be categorized into two main approaches: the geostatistical approach, which uses geographic coordinates (latitude and longitude), and the lattice approach, which considers the relative position of an area to neighboring regions and our method is able to deal with both the framework. A comprehensive literature review, including Bayesian approaches, can be found in Hanson and Zhou (2014). Our interest lies in a frequentist approach to the spatial Proportional Hazards model, for which various models and methods are now available to facilitate inference. Examples include the semiparametric estimating equations method proposed by Li and Lin (2006) and the composite likelihood method Paik and Ying (2012). The main issue with these proposals is that the dependence structure does not consider the shape of the domain but only the distance among the locations. This can be problematic when the boundary of the domain plays a crucial role in the spatial dependence. In this paper, we propose a novel approach based on recent developments in differential regularization, as discussed in Wilhelm and Sangalli (2016). One of the main features of this approach is to be able to deal with domains having complex shapes, characterized for instance by strong concavities or holes, and where the shape of the domain influences the behavior of the phenomenon.
@@
We demonstrate that the methodology introduced in Wilhelm and Sangalli (2016), which concerns a generalized regression model for spatially distributed data when the response variable has a distribution within the exponential family, can be extended to a spatial proportional hazard model with randomly right-censored data. Specifically, we optimize a penalized log-likelihood function that includes a roughness penalty term based on a differential quantity of the spatial field, calculated over the domain of interest. This approach utilizes the so called Generalized Spatial Regression with PDE penalization (GSR-PDE) method (Wilhelm and Sangalli, 2016). Our proposed models leverage finite element methods on a triangulation of the domain to achieve precise estimates of the spatial field, particularly in domains with complex geometries, where triangulation effectively captures the intricacies of the area.
@@ 2 Model
We consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with a boundary $\partial\Omega$ that is regular and belongs to $C^{2}(\mathbb{R}^{2})$. Let $p_{1},\ldots,p_{n}\in\Omega$ denote $n$ points, where $p_{i}=(p_{1i},p_{2i})$. Each point $p_{i}$ is associated with a time-to-event random variable $T_{i}$. Due to censoring, $T_{i}$ cannot be observed directly; instead, we obtain a realization of $(Y_{i},\Delta_{i})$, denoted by $(y_{i},\delta_{i})$, where $Y_{i}=\min(T_{i},C_{i})$, $\Delta_{i}=I(T_{i}\leq C_{i})$, $C_{i}$ is a censoring random variable, and $I(\cdot)$ is the indicator function. Additionally, we observe a covariate vector $x_{i}\in\mathbb{R}^{q}$ at each location $p_{i}$. We assume that $T_{i}$ is independent of $C_{i}$ conditional on $x_{i}$. We use a proportional hazard model for the random variables $\{T_{i}\}_{i=1}^{n}$. Specifically, let $\lambda_{i}(t\mid x)$ denote the conditional hazard function of $T_{i}$ given $x_{i}=x$. We assume that
$\lambda_{i}(t\mid x)=\lambda_{0}(t)\exp(x^{\top}\beta+f(p_{i})),$ (1)
where $\lambda_{0}$ is the unspecified baseline hazard, $\beta\in\mathbb{R}^{q}$ is a vector of regression coefficients, and $f$ is a smooth function over $\Omega$ belonging to a suitable functional space $\mathcal{F}$. Note that $f$ is identifiable up to an additive constant, so we assume that $\mathcal{F}$ contains only functions such that at a given point $p^{*}\in\Omega$ it holds $f(p^{*})=0$.
We propose to estimate the regression coefficients $\beta\in\mathbb{R}^{q}$ and the spatial field $f\in\mathcal{F}$ by maximizing a penalized log-partial-likelihood functional:
@@
where $\mathcal{L}(\beta,f)$ is the standard partial likelihood defined as
@@
We show that maximizing the penalized log-likelihood function (2) is equivalent to minimizing the penalized least-squares functional considered in Sangalli et al. (2013). The quadratic form of this functional allows for an analytical characterization of its minimum (or equivalently, the maximum of the penalized log-likelihood functional), and thus for the estimation of $\hat{\beta}\in\mathbb{R}^{q}$ and $\hat{f}\in\mathcal{F}$. As in Wilhelm and Sangalli (2016), which considers an exponential family, we cannot directly apply the standard PIRLS algorithm developed by O’sullivan et al. (1986) for maximizing a penalized log-likelihood functional in generalized additive models. This is because the penalized log-likelihood in (2) involves a function parameter, the spatial field $f$, and the maximization occurs over the space $\mathbb{R}^{q}\times\mathcal{F}$, where $\mathcal{F}$ is infinite-dimensional. Instead we recall the functional version of the PIRLS algorithm proposed in Wilhelm and Sangalli (2016) to find an approximate solution to the estimation problem.
@@ 3 Functional version of the PIRLS algorithm
$\mathcal{L}_{p}(\beta,f)=\mathcal{L}(\beta,f)-\frac{\mu}{2}m(f,f),$ (4)
@@
$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
@@
$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f),$
$l_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)},$
$v_{i}(\beta,f)$ $=\theta_{i}(\beta,f)/l_{i}(\beta,f).$
First, we show that the problem of maximizing 4 is equivalent to minimizing the following functional $\mathcal{J}_{\mu}(\beta,f)$ with respect to $(\beta,f)$:
$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$, where such quantities are supposed to be constant. As, in fact, $l$ and $V$ also depend on $(\beta,f)$, this suggests an iterative scheme for the solution of the estimation problem. Let $(\beta^{(k)},f_{n}^{(k)})$ be the value of $(\beta,f)$ after $k$ iterations of the algorithm, and let $\theta^{(k)}$ the estimate of the vector $\theta$ after $k$ iterations, and define $\Theta^{(k)}=\text{diag}(\theta^{(k)})$. Define also $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X^{\top}\beta^{(k)}+f^{(k)}$, and let ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$. The first order approximation of $\theta(\beta,f)$ in the neighborhood of the current value of $\theta^{(k)}$ has to be considered in the space $\mathbb{R}^{q}\times\mathcal{F}$ and yields to the following quadratic approximation of $\mathcal{J}_{\mu}(\beta,f)$:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X^{\top}\beta-f)\right\|^{2}+\mu m(f,f).$ (5)
@@
    Compute $z^{(k)}$ and $W^{(k)}$ based on $(\beta^{(k)},f_{n}^{(k)})$.
@@
    Next, find $\beta$ and $f_{n}$ that minimize (5).
@@
    Perform the standard proportinal hazards model estimation using $X$ and $f_{n}^{(k)}$ as covariate and denote by $\hat{\beta}$ and $\hat{c}$ the resulting regression coeffiecint for $X$ and $f_{n}^{(k)}$, respectively.
@@
    Define $\beta^{(k+1)}=\hat{\beta}$ and and $f_{n}^{(k+1)}=\hat{c}(f_{n}^{(k)}-f_{n}^{(k)}(p^{*}))$.
@@ 4 Numerical Implementation using the Finite Element Method
$\tilde{J}_{\lambda}(\beta,f)=\|W^{1/2}(z-X\beta-f_{n})\|^{2}+\lambda\int_{\Omega}(\Delta f)^{2}.$ (6)
For simplicity, we omit the dependence on the iteration counter $k$ in this notation and throughout the following sections. Let us now consider what type of space $\mathcal{F}$ is suitable for the current problem. To do so, we introduce the Sobolev space $H^{m}(\Omega)$, which is a Hilbert space comprising all functions in $L^{2}(\Omega)$ along with all their distributional derivatives up to order $m$. Since the roughness penalty term $\int_{\Omega}(\Delta f)^{2}$ must be well-defined, it follows that $\mathcal{F}\subset H^{2}(\Omega)$. According to the Sobolev embedding theorem, $H^{2}(\Omega)\subset C^{0}(\Omega)$. Therefore, any function $f\in H^{2}(\Omega)$ is continuous and can be evaluated pointwise, allowing us to compute the vector $f_{n}$ in the partial likelihood term. To ensure the uniqueness of the minimizer in (6), appropriate boundary conditions must be imposed. These conditions specify the desired behavior of the estimated function $f$ at the boundaries of the domain of interest. Typically, constraints can be set on the value of $f$ at the boundary $\partial\Omega$, i.e., $f|_{\partial\Omega}=\gamma_{D}$ (Dirichlet boundary conditions), or on the flux of the function through the boundary, i.e., $\partial_{n}f|_{\partial\Omega}=(\nabla f)^{\top}\mathbf{n}=\gamma_{N}$ (Neumann boundary conditions), where $\mathbf{n}$ is the outward-pointing normal unit vector to the boundary, and $\nabla f=(\partial f/\partial p_{1},\partial f/\partial p_{2})^{\top}$ represents the gradient of $f$. If $\gamma_{D}$ or $\gamma_{N}$ are zero functions, the boundary conditions are termed homogeneous. It is also possible to apply different boundary conditions to various parts of the boundary, forming a partition of $\partial\Omega$. To ensure the uniqueness of the minimization problem (4.1), we consider the space:
@@
For more details on general boundary conditions, see Azzimonti et al. (2014).

### v2

@@
Abstract
This paper presents a semiparametric proportional hazards model designed to handle spatially varying covariate functions, applicable to both geostatistical and areal data observed on irregular spatial domains. The model is estimated through the maximization of a penalized partial likelihood, with a roughness penalty term based on a differential of the spatial field over the target domain. The finite element method is employed for efficient estimation, enabling a piecewise polynomial surface representation of the spatial field. We apply this method to analyze response time data from the London Fire Brigade.
Keywords: spatial proportional hazards model, differential regularization, semiparametric model, geostatistical data, areal data, finite element method, spatial field, surivival analysis
@@ 1 Introduction
Survival analysis has long been an essential tool across various fields, including medicine, engineering, and social sciences, for analyzing time-to-event data. In this paper, we focus on one of the most widely used models within survival analysis: the Proportional Hazards (PH) model, first introduced by Cox, (1972). The PH model assesses how covariates influence the hazard, or risk, of an event happening at any given moment in time, relative to a baseline risk. This model operates under the assumption that the effect of each covariate is proportional and constant over time, meaning that covariates alter the hazard rate by a multiplicative factor, without assuming a specific underlying time distribution for the event.
@@
In recent years, survival analysis has advanced significantly in modeling various dependencies within survival data. Much research has focused on the relationship between censoring and time variables Deresa and Keilegom, (2024), a critical aspect when handling censored data. However, other forms of dependence have received comparatively less attention. This paper addresses a scenario where the occurrence of an event is linked to its geographic location, thereby introducing spatial dependence into the analysis. Specifically, the probability of an event may be influenced by spatial characteristics associated with the event’s location. We call the Proportional Hazards model that incorporates these spatial factors as the Spatial Proportional Hazards (SPH) model.
Spatial factors can generally be categorized into two main approaches: the geostatistical approach, which uses geographic coordinates (latitude and longitude), and the lattice approach, which considers the relative position of an area to neighboring regions, and our method is able to deal with both the framework.
A comprehensive literature review, including Bayesian approaches, is available in Hanson and Zhou, (2014). In this work, however, we focus on a frequentist approach to the SPH model, for which various inference methods have been developed. Examples include the semiparametric estimating equations method by Li and Lin, (2006) and the composite likelihood method by Paik and Ying, (2012). A significant limitation of these methods is their reliance on a dependence structure that accounts only for the distance between locations, ignoring the domain’s shape. This approach can lead to inaccuracies, especially when spatial dependence is influenced by the domain boundary.
Additionally, most semiparametric SPH models in the literature use a spline estimator to parameterize the baseline hazard function as infinite-dimensional parameters, rather than leaving this component fully unspecified, as in Cox’s traditional partial likelihood approach (Cox,, 1972). This additional parameterization requires estimating more parameters, leading to reduced robustness and accuracy in the estimation process.
In contrast, our method is grounded in the traditional partial likelihood approach, focusing specifically on the key multiplicative component of the model. Unlike other methodologies that introduce spatial dependence through a frailty-based approach—where the spatial domain is clustered and each cluster has a distinct effect on the hazard following a specified distribution—we incorporate a non-parametric spatial field defined over the entire domain to capture spatial dependence in the data.
For the estimation, we propose a novel method that leverages recent advancements in differential regularization, as discussed by Wilhelm and Sangalli, (2016). A key advantage of our approach is its capacity to accommodate domains with complex shapes, such as those characterized by strong concavities or holes, where the domain geometry significantly impacts the behavior of the observed phenomenon.
@@
We demonstrate that the methodology introduced in Wilhelm and Sangalli, (2016), which concerns a generalized regression model for spatially distributed data when the response variable has a distribution within the exponential family, can be extended to a spatial proportional hazard model with randomly right-censored data. Specifically, we optimize a penalized log-likelihood function that includes a roughness penalty term based on a differential quantity of the spatial field, calculated over the domain of interest. This approach utilizes the so called Generalized Spatial Regression with PDE penalization (GSR-PDE) method (Wilhelm and Sangalli,, 2016). Our proposed models leverage finite element methods on a triangulation of the domain to achieve precise estimates of the spatial field.
@@ 2 Model
We consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with a boundary $\partial\Omega$ that is regular and belongs to $C^{2}(\mathbb{R}^{2})$. Let $\bm{p}_{1},\ldots,\bm{p}_{n}\in\Omega$ denote $n$ points, where $\bm{p}_{i}=(p_{1i},p_{2i})$. Each point $\bm{p}_{i}$ is associated with a time-to-event random variable $T_{i}$. Due to censoring, $T_{i}$ cannot be observed directly; instead, we obtain a realization of $(Y_{i},\Delta_{i})$, denoted by $(y_{i},\delta_{i})$, where $Y_{i}=\min(T_{i},C_{i})$, $\Delta_{i}=I(T_{i}\leq C_{i})$, $C_{i}$ is a censoring random variable, and $I(\cdot)$ is the indicator function. Additionally, we observe a covariate vector $\bm{x}_{i}\in\mathbb{R}^{q}$ at each location $p_{i}$. We assume that $T_{i}$ is independent of $C_{i}$ conditional on $\bm{x}_{i}$. We use a proportional hazard model for the random variables $\{T_{i}\}_{i=1}^{n}$. Specifically, let $\lambda_{i}(t\mid\bm{x})$ denote the conditional hazard function of $T_{i}$ given $\bm{x}_{i}=\bm{x}$. We assume that
$\lambda_{i}(t\mid x)=\lambda_{0}(t)\exp(\bm{x}^{\top}\bm{\beta}+f(\bm{p}_{i})),$ (1)
where $\lambda_{0}$ is the unspecified baseline hazard, $\beta\in\mathbb{R}^{q}$ is a vector of regression coefficients, and $f$ is a smooth function from $\Omega$ to $\mathbb{R}$ belonging to a suitable functional space $\mathcal{F}$. Thus, we will interpret the model in terms of hazards: the vector of coefficients $\bm{\beta}$ quantifies the influence of the covariates $\bm{x}_{i}$ on the hazard function, so each component of $\beta$ can be interpreted as the log hazard ratio associated with a one-unit increase in the corresponding covariate, holding other factors constant. The function $f(\bm{p}_{i})$, instead, captures the spatial variation in hazards that is not accounted for by the covariates. Specifically, $f(\bm{p}_{i})$ allows for the identification of localized risk factors affecting the hazard at different points in the domain, revealing areas with higher or lower risks relative to the baseline hazard $\lambda_{0}(t)$. Thus, the interpretation of hazards is twofold: it incorporates covariate effects via $\bm{\beta}$ and captures residual spatial variation through $f(\bm{p})$, providing a comprehensive view of risk that addresses both individual and spatial heterogeneity.
Note that $f$ is identifiable up to an additive constant, so we assume that $\mathcal{F}$ contains only functions such that at a given point $\bm{p}^{*}\in\Omega$ it holds $f(\bm{p}^{*})=0$. Note that alternative identifiability conditions can be considered, such as $\int_{\Omega}f=1$. We propose to estimate the regression coefficients $\bm{\beta}\in\mathbb{R}^{q}$ and the spatial field $f\in\mathcal{F}$ by maximizing a penalized log-partial-likelihood functional:
@@
where $\mathcal{L}(\bm{\beta},f)$ is the standard log-partial-likelihood for the PH model, defined as
@@
We show that maximizing the penalized log-likelihood function (2) is equivalent to minimizing the penalized least-squares functional considered in Sangalli et al., (2013). The quadratic form of this functional allows for an analytical characterization of its minimum (or equivalently, the maximum of the penalized log-likelihood functional), and thus for the estimation of $\hat{\bm{\beta}}\in\mathbb{R}^{q}$ and $\hat{f}\in\mathcal{F}$. As in Wilhelm and Sangalli, (2016), which considers an exponential family, we cannot directly apply the standard PIRLS algorithm developed by O’sullivan et al., (1986) for maximizing a penalized log-likelihood functional in generalized additive models. This is because the penalized log-likelihood in (2) involves a function parameter, the spatial field $f$, and the maximization occurs over the space $\mathbb{R}^{q}\times\mathcal{F}$, where $\mathcal{F}$ is infinite-dimensional. Instead we recall the functional version of the PIRLS algorithm proposed in Wilhelm and Sangalli, (2016) to find an approximate solution to the estimation problem.
@@ 3 Functional version of the PIRLS algorithm
$\mathcal{L}_{p}(\bm{\beta},f)=\mathcal{L}(\bm{\beta},f)-\frac{\mu}{2}m(f,f),$ (4)
@@
$\theta_{i}(\bm{\beta},f)$ $=\exp(\bm{x}_{i}^{\top}\bm{\beta}+f(p_{i})),$
@@
$s_{i}(\bm{\beta},f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\bm{\beta},f),$
$l_{i}(\bm{\beta},f)$ $=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\bm{\beta},f)},$
$v_{i}(\bm{\beta},f)$ $=\theta_{i}(\bm{\beta},f)/l_{i}(\bm{\beta},f).$
First, we show that the problem of maximizing (4) is equivalent to minimizing the following functional $\mathcal{J}_{\mu}(\beta,f)$ with respect to $(\beta,f)$:
$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\bm{\delta}/\bm{l}-\theta)\|^{2}+\mu m(f,f),$
where $\bm{\delta}=(\delta_{1},\dots,\delta_{n})^{\top}$, $\bm{l}=(l_{1},\dots,l_{n})^{\top}$, $\bm{\delta}/\bm{l}=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$, and $V=\text{diag}(v_{1},\dots,v_{n})$, where such quantities are supposed to be constant. As, in fact, $\bm{l}$ and $V$ also depend on $(\bm{\beta},f)$, this suggests an iterative scheme for the solution of the estimation problem. Let $(\bm{\beta}^{(k)},f_{n}^{(k)})$ be the value of $(\bm{\beta},f)$ after $k$ iterations of the algorithm, and let $\bm{\theta}^{(k)}$ the estimate of the vector $\bm{\theta}=(\theta_{1},\dots,\theta_{n})^{\top}$ after $k$ iterations, and define $\Theta^{(k)}=\text{diag}(\bm{\theta}^{(k)})$. Let $X$ be the $q\times n$ matrix where the column $i$ is equal to $x_{i}$ and define also $\bm{z}^{(k)}=(\Theta^{(k)})^{-1}(\bm{\delta}/\bm{l}-\bm{\theta}^{(k)})+X^{\top}\bm{\beta}^{(k)}+f^{(k)}$, and let ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\bm{\beta}^{(k)},f^{(k)})\theta_{i}^{2}(\bm{\beta}^{(k)},f^{(k)})$. The first order approximation of $\theta(\bm{\beta},f)$ in the neighborhood of the current value of $\bm{\theta}^{(k)}$ has to be considered in the space $\mathbb{R}^{q}\times\mathcal{F}$ and yields to the following quadratic approximation of $\mathcal{J}_{\mu}(\beta,f)$:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\bm{\beta},f)=\left\|({W}^{(k)})^{1/2}(\bm{z}^{(k)}-X^{\top}\bm{\beta}-f)\right\|^{2}+\mu m(f,f).$ (5)
@@
    Compute $\bm{z}^{(k)}$ and $W^{(k)}$ based on $(\bm{\beta}^{(k)},f_{n}^{(k)})$.
@@
    Next, find $\bm{\beta}$ and $f_{n}$ that minimize (5).
@@
    Perform the standard proportinal hazards model estimation using $X$ and $f_{n}^{(k)}$ as covariate and denote by $\hat{\bm{\beta}}$ and $\hat{c}$ the resulting regression coeffiecint for $X$ and $f_{n}^{(k)}$, respectively.
@@
    Define $\bm{\beta}^{(k+1)}=\hat{\bm{\beta}}$ and and $f_{n}^{(k+1)}=\hat{c}(f_{n}^{(k)}-f_{n}^{(k)}(p^{*}))$.
@@ 4 Numerical Implementation using the Finite Element Method
$\tilde{J}_{\lambda}(\bm{\beta},f)=\|W^{1/2}(z-X^{\top}\bm{\beta}-f_{n})\|^{2}+\lambda\int_{\Omega}(\Delta f)^{2}.$ (6)
For simplicity, we omit the dependence on the iteration counter $k$ in this notation and throughout the following sections. Let us now consider what type of space $\mathcal{F}$ is suitable for the current problem. To do so, we introduce the Sobolev space $H^{m}(\Omega)$, which is a Hilbert space comprising all functions in $L^{2}(\Omega)$ along with all their distributional derivatives up to order $m$. Since the roughness penalty term $\int_{\Omega}(\Delta f)^{2}$ must be well-defined, it follows that $\mathcal{F}\subset H^{2}(\Omega)$. According to the Sobolev embedding theorem, $H^{2}(\Omega)\subset C^{0}(\Omega)$. Therefore, any function $f\in H^{2}(\Omega)$ is continuous and can be evaluated pointwise, allowing us to compute the vector $f_{n}$ in the partial likelihood term. To ensure the uniqueness of the minimizer in (6), appropriate boundary conditions must be imposed. These conditions specify the desired behavior of the estimated function $f$ at the boundaries of the domain of interest. Typically, constraints can be set on the value of $f$ at the boundary $\partial\Omega$, i.e., $f|_{\partial\Omega}=\gamma_{D}$ (Dirichlet boundary conditions), or on the flux of the function through the boundary, i.e., $\partial_{n}f|_{\partial\Omega}=(\nabla f)^{\top}\mathbf{n}=\gamma_{N}$ (Neumann boundary conditions), where $\mathbf{n}$ is the outward-pointing normal unit vector to the boundary, and $\nabla f=(\partial f/\partial p_{1},\partial f/\partial p_{2})^{\top}$ represents the gradient of $f$. If $\gamma_{D}$ or $\gamma_{N}$ are zero functions, the boundary conditions are said to be homogeneous. It is also possible to apply different boundary conditions to various parts of the boundary, forming a partition of $\partial\Omega$. To ensure the uniqueness of the minimization problem (4.1), we consider the space:
@@
For more details on general boundary conditions, see Azzimonti et al., (2014).